# Generalization Below the Edge of Stability: The Role of Data Geometry

### v1

@@ Appendix I Flat Interpolating Two-Layer ReLU Networks on the Unit Sphere
By (88) and (89), the model interpolates on nonzero labels because $f({\bm{x}}_{k})=a_{k}(1-b_{k})=y_{k}$ for $k\in I_{\neq 0}$, and it also interpolates zero labels since all constructed units are inactive on ${\bm{x}}_{i}$ when $i\notin J_{\neq 0}$, hence $f({\bm{x}}_{i})=0=y_{i}$.
@@
$\tilde{v}_{k}:=\operatorname{sign}(v_{k})\in\{\pm 1\},\qquad\tilde{{\bm{w}}}_{k}:=|v_{k}|\,{\bm{w}}_{k},\qquad\tilde{b}_{k}:=|a_{k}|\,b_{k}.$ (90)
@@
so interpolation is preserved. Moreover, the activation pattern on the dataset is unchanged because (88) has strict inequalities and $|a_{i}|>0$. At ${\bm{x}}_{i}$ we have the (post-rescaling) pre-activation
$\tilde{z}_{k}\;:=\;\tilde{{\bm{w}}}_{k}^{\mathsf{T}}{\bm{x}}_{k}-\tilde{b}_{k}=|a_{k}|\,(1-b_{k})=|y_{k}|\;>\;0,\qquad|\tilde{v}_{i}|=1.$ (92)
@@
$\frac{\partial f({\bm{x}}_{i})}{\partial b_{k}}$ $=\mathds{1}\{{\bm{w}}_{k}^{\mathsf{T}}{\bm{x}}>b_{k}\}\,v_{k},\quad$ $\frac{\partial f({\bm{x}}_{i})}{\partial\beta}$ $=1.$
@@
If we remove the output bias term $\beta$ from the parameters, then the bottom row of 100 will be remove and thus term $\sum_{i}u_{i}$ in (102) will be removed. ∎

### v2

@@ Appendix I Flat Interpolating Two-Layer ReLU Networks on the Unit Sphere
By (92) and (93), the model interpolates on nonzero labels because $f({\bm{x}}_{k})=v_{k}(1-b_{k})=y_{k}$ for $k\in I_{\neq 0}$, and it also interpolates zero labels since all constructed units are inactive on ${\bm{x}}_{i}$ when $i\notin I_{\neq 0}$, hence $f({\bm{x}}_{i})=0=y_{i}$.
@@
$\tilde{v}_{k}:=\operatorname{sign}(v_{k})\in\{\pm 1\},\qquad\tilde{{\bm{w}}}_{k}:=|v_{k}|\,{\bm{w}}_{k},\qquad\tilde{b}_{k}:=|v_{k}|\,b_{k}.$ (94)
@@
so interpolation is preserved. Moreover, the activation pattern on the dataset is unchanged because (92) has strict inequalities and $|v_{k}|>0$. At ${\bm{x}}_{i}$ we have the (post-rescaling) pre-activation
$\tilde{z}_{k}\;:=\;\tilde{{\bm{w}}}_{k}^{\mathsf{T}}{\bm{x}}_{k}-\tilde{b}_{k}=|v_{k}|\,(1-b_{k})=|y_{k}|\;>\;0,\qquad|\tilde{v}_{k}|=1.$ (96)
@@
$\frac{\partial f({\bm{x}}_{i})}{\partial b_{k}}$ $=\mathds{1}\{{\bm{w}}_{k}^{\mathsf{T}}{\bm{x}}>b_{k}\}\,-v_{k},\quad$ $\frac{\partial f({\bm{x}}_{i})}{\partial\beta}$ $=1.$
@@
If we remove the output bias term $\beta$ from the parameters, then the bottom row of 104 will be remove and thus term $\sum_{i}u_{i}$ in (106) will be removed. ∎